The cost in dollars of manufacturing w wing nuts is given by the expression 0.07w + 13.33. Find the cost when 375 wing nuts are made. What is the average cost to manufacture one wing nut?
Answer by checkley75(3666)

.07W+13.33
.07*375+13.33=26.25+13.33=39.58 COST TO MAKE 375 WING NUTS.
39.58/375=.1055 OR $.11 PER WING NUT.